OpenStudy (anonymous):

Ten is the geometric mean of 4 and what number? 2 40 25 8

OpenStudy (tpaulus):

\[10 = \sqrt{x * 4}\] Solve for x, by first squaring... Don't forget to check for extraneous solutions

OpenStudy (anonymous):

So 40?

jhonyy9 (jhonyy9):

squared both sides and will get 100 = 4x divide both sides by 4 so x = 100/4 = ???

OpenStudy (anonymous):

25